# Convergence on the Proton Drip-Line in Thulium

### v3

@@ IV Conclusion and Outlook
Our update to the literature masses also highlights the importance of measuring and reporting not only the masses of previously unmeasured isotopes but also confirmations of literature mass values which provide input for various models and predictions. Heavier exotic species pose challenges that arise from the abundance of isomers and isobaric molecules that must be identified and, therefore, must be treated with caution. The mass values presented here provide a new benchmark for the models used to predict the binding energies of unmeasured nuclei in the neutron-deficient region near N=82 and establish the Z=69 proton drip-line.
@@ V Acknowledgements
This work was further supported by the German Research Foundation (DFG), Grant No. 422761894; by the German Federal Ministry for Education and Research (BMBF), Grants No. 05P16RGFN1 and No. 05P19RGFN1; by the Hessian Ministry for Science and Art through the LOEWE Center HIC for FAIR; by the JLU and GSI Helmholtzzentrum für Schwerionenforschung under the JLU-GSI strategic Helmholtz partnership agreement.
For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence to any author accepted manuscript version arising.

### v5

@@ IV Conclusion and Outlook
Our update to the literature masses also highlights the importance of measuring and reporting not only the masses of previously unmeasured isotopes but also confirmations of literature mass values which provide input for various models and predictions. Heavier exotic species pose challenges that arise from the abundance of isomers and isobaric molecules that must be identified and, therefore, must be treated with caution. The mass values presented here provide a new benchmark for the models used to predict the binding energies of unmeasured nuclei in the neutron-deficient region near and establish the proton drip-line.
@@ V Acknowledgements
This work was further supported by the German Research Foundation (DFG), Grant No. 422761894; by the German Federal Ministry for Education and Research (BMBF), Grants No. 05P16RGFN1 and No. 05P19RGFN1; by the Hessian Ministry for Science and Art through the LOEWE Center HIC for FAIR; and by the JLU and GSI Helmholtzzentrum für Schwerionenforschung under the JLU-GSI strategic Helmholtz partnership agreement.